A lottery has a very large number of tickets, one in every 500 at which entitles the purchase to prize. Calculate the minimum number of tickets the agent must sell to have 90% chance of selling at least one prize winning ticket.


1
Expert's answer
2021-03-31T16:54:48-0400

The probability that ticket entitle the purchaser to prize is

"p=\dfrac{1}{500}=0.002"

If purchaser buy N tickets, the probability that none of them will be winning according to the formula of probability of independent events intersection:


"P(lose)=(1-p)^N"

So, the probability that purchaser has at least one prize winning ticket is:


"P(at\ least \ one\ prize)=1-P(lose)"

"=1-(1-p)^N"

By condition, this value should be not less than

"p_0=0.9"

"1-(1-p)^N\geq p_0"

"1-(1-0.002)^N\geq 0.9"

"(0.998)^N\leq 0.1"

"N\ln(0.998)\leq\ln(0.1)"

"\ln(0.998)<0: N\geq\dfrac{\ln(0.1)}{\ln(0.998)}"

"N\geq1151"

So, the minimum no. of tickets the agent must sell to have 90% chance of selling at least one prize winning ticket is 1151.